# Beyond the Dashboard: Investigating Distracted Driver Communication Preferences for ADAS

### v1

@@ IV-C Adaptive Communication Mode Questionnaire
To test H3, participants were asked if they would feel safer in a car with such a system over a car without such a system as both a driver and a passenger. We posit that drivers’ perceived sense of safety indicates a preference for a system as we assume that most rational drivers value safety. The dependent variable was how safe participants felt, and the independent variables were the demographic details of the participants [^3] Discussed in Appendix -C..
Implementing an Adaptive Communication Module would require in-car driver monitoring to assess drivers current distraction state. Currently, gaze based distraction estimation is a state-of-the-art non-invasive method to estimate driver distraction which only requires eye gaze data through the use of ubiquitous low-cost RGBD cameras [2, 3, 38]. While, other methods of estimating driver distraction such as EEGs and heart rate monitors are more accurate and have been used in previous naturalistic user studies, they are more invasive and require additional complicated hardware [39, 40, 5]. Additionally, the general population is also unfamiliar with these devices and their inclusion would require additional context which would make the survey more complex and arduous. Therefore, participants were asked if they would be comfortable with their audio and facial videos being recorded by their vehicle. The dependent variable was how comfortable participants felt, and the independent variables were the demographic details of the participants. We verify this claim by asking participants about their distractions in the driving experience questionnaire [^4] Results are presented in the Appendix -D.. We posit that the user’s comfort with a plausible detection system in addition to their perceived sense of safety indicates their trust in such a system to test H4.
IV-D Participants
Of the 147 participants, 90 participants were recruited through Amazon MTurk, while the rest were recruited through the other channels mentioned Section IV-A. Three participants were excluded from the study as they did not meet the inclusion criteria. Therefore, a total of 144 participants (mean age=37; median age=39; range=18-65; 87 males; 57 females) were considered in the analysis presented in Section V.
@@ V Results and Discussion
In this section, we analyze the responses obtained from our user study. We present qualitative and quantitative results to support our hypotheses. We also derive conclusions regarding the usage of modes based on the driving scenario. Additionally, we recommend design changes for the communication paradigms between drivers and ADAS.
V-A The relation between the modes and characteristics of a driving scenario
@@
Figure 3: The distribution of the modes of communication based on the hazardous nature of the driving scenario (left to right in decreasing order), regardless of distraction status. Each section of the vertical bars shows the number of times the mode was picked as the first (bottom), second (middle), and third (top) choice respectively.
Multiple statistical tests were performed to compare the effect of the different driving scenarios (independent variable) on the ranks assigned to the modes (dependent variable): (1) A one-way ANOVA revealed that there was a statistically significant difference between the mode rankings for the five driving scenarios with F-score of $28.19$ with $p<0.05$ on average for all modes; (2) A Chi-square Goodness of Fit test was performed to determine whether the weighted mode selection frequencies (calculated as $3\times f_{\text{rank 1}}+2\times f_{\text{rank 2}}+1\times f_{\text{rank 3}}$, where $f_{\text{rank $i$}}$ is the frequency of the mode selected at rank $i$) was equal between the different driving scenarios. The test revealed that the mode frequencies differ by the driving scenario with $\chi^{2}(3,5)=158.85,\ p<0.001$ on average for all modes. Other Chi-square tests that were used to determine whether the frequencies of the modes at each rank was different between each driving scenario also indicated similar results. Both statistics discussed above indicate a statistically significant difference in mode frequencies between the different driving scenarios.
@@
As seen in Figure 3, participants prefer different modes of communication depending on the hazardous nature of the driving scenario. Particularly, participants prefer direct modes that can be interpreted quickly such as repeated sound alerts and vibrations for emergent, dangerous scenarios. For low-risk scenarios, participants prefer modes that serve as reminders and can be interpreted without haste, such as a singular sound beep or spoken language alerts. These conclusions are in line with the guidelines presented by Campbell et al.  for HVI as well as the results presented by Du et al[10, 26].
@@
Meanwhile, some modes such as non-flashing icons and text or images on the dashboard are not preferred by participants in any setting. These modes are consistently the least selected modes across all driving scenarios presented in the study. Notably, these modes are usually reserved for long term information regarding vehicle health and might be unsuitable for conveying information for the driving scenarios depicted in the study. However, a growing number of vehicles implement text and image alerts on the dashboard and central consoles as common modes for all types of automotive information [41]. We suggest that these modes should be used sparsely and with careful consideration.
@@
The difference between the frequency of the modes for the most and least hazardous scenario: Pedestrian collisions and Standing still at a Green light, clearly indicates that drivers prefer different modes of communication based on the risk involved in a scenario. This conclusion is also reaffirmed by the steady change of the frequency of modes as the risk in the scenario decreases, as seen from left to right in Figure 3. However, we note that the final rankings of the different modes is consistent across the first four driving scenarios. While this consistency indicates an overall unchanging sentiment, we note that the deviations in the mode frequencies for each scenario are significant as proven by the one-way ANOVA test. Moreover, there are significant deviations in the frequency of selections for each mode that imply the importance of personalization for an adaptive communication framework. Therefore, we conclude that H1 is true. We also note that our findings are consistent with that of Du et al[26]. This confirmation of H1 provides credence to the results presented below as H1 was intended as a confirmation hypothesis.
V-B Evolution of driver preferences with changing attention
@@
The responses from the video questionnaires can be treated as a ranking of the top-3 preferences out of the 10 choices presented in Table II. These top-3 rankings can be transformed to fully ranked lists of all the modes by allowing for ties. We choose to assign all modes that were not selected by users a tied rank of 4. As mentioned in Section IV-B, we posit that H2 holds true if the ranked lists for the non-augmented and augmented versions of the videos are not fully correlated. We use the following rank correlation metrics to compare the responses for each video pair:
  1. 1.
@@
    Kendall’s $\tau$ [42]: The difference between the fraction of ranking pairs that appear in order and the fraction of ranking pairs that appear out of order. Values close to 1 indicate strong agreement while values close to -1 indicate strong disagreement.
  2. 2.
@@
    Spearman’s $\rho$ [43]: The Pearson correlation coefficient between the rank variables. Values close to 1 indicate agreement while values close to -1 indicate disagreement.
  3. 3.
@@
    Rank Biased Overlap (RBO) [44]: The expected average overlap between the two rankings at incremental overlap depths. Values close to 1 indicate agreement, while values close to 0 indicates disagreement.
@@
The different metrics measure different types of correlation between the two rankings. Both Kendall’s $\tau$ and Spearman’s $\rho$ are unweighted measures and only check to see the ordering of the two rankings is the same without placing any weight on the change in ranking (i.e. scores for items with rank 1 and 2 would be the same as the score for items with rank 2 and 3). Whereas, the RBO metric assigns significance to the actual rank changes and thus weighs the changes in ranking more than just their order. Therefore, RBO is more sensitive to changes in ranking. All analysis was carried out in python using the scipy and rbo packages [45, 46].
@@
TABLE III: The Correlations Between the Ranked Modes of Communication for the Non-augmented and Augmented Videos. The p-values for all metrics was less than 0.0004.
@@
Table III shows the correlations for these metrics aggregated according to the driving scenario. The correlations were aggregated as simple averages and the p-values were aggregated using the Fisher exact test [47]. All metrics show that the rankings for the non-augmented and augmented versions are not fully correlated but only slightly positively correlated. We speculate that this result is a consequence of the modes of communication being more closely related to the critical nature of the driving scenario than the current distraction state. We further conjecture that the positive correlation implies that while driver preferences for the mode of communication change, drivers are not completely changing their minds on what modes best suit the current driving scenario. We also observe that the correlation between the rankings decreases as the scenarios become less hazardous. The correlation values ranging from 0.5 to 0.7 reinforces this belief. Additionally, we believe that these values are also a result of converting the partially ranked top-3 list to a fully ranked list with ties at rank 4. This policy ensures that a large portion of the ranking lists are the same, since 7 out of 10 options were all ranked at the same rank. This procedure heavily biases the correlation observed between the rankings to be positive. Based on these findings we conclude that H2 is true.
V-C Sentiments towards an Adaptive Communication Module
@@
Figure 4: The distribution of participant responses for perceived safety in a vehicle with an Adaptive Communication Module as a driver (top) and passenger (bottom).
@@
Figure 5: The distribution of participant responses for their comfort with audio (top) and visual (bottom) recordings in their vehicles.
@@
The responses for participants perceived safety in a vehicle with an adaptive communication module are shown in Figure 4. The mean score (1=Strongly disagree; 5=Strongly agree) for perceived safety as a driver and passenger are $3.93\pm 0.91$ and $3.95\pm 1.03$, respectively. Over 80% and 77% of respondents believe that they would feel safer in a vehicle with AdaCoM as a driver and as a passenger, respectively. No bias between demographic variables for the participants and their perceived safety was indicated by model fitting using an Ordinal Multinomial Probit model. Therefore, we conclude that H3 holds true. However, we note that this finding is based on responses obtained from asking participants about a hypothetical system. While the results presented here motivate the need for development and testing for such a system, there is no strong conclusion regarding its reception.
@@
The responses for the participants comfort with a plausible implementation of an AdaCoM system are shown in Figure 5. The mean score for comfort with audio and video recordings are $2.32\pm 1.20$ and $2.14\pm 1.19$, respectively. Over 60% and 67% of respondents are uncomfortable having their audio and video recorded by their vehicles, respectively. Model fitting using an Ordinal Multinomial Probit model showed that the age of the participants was a statistically significant $(p=0.056)$ indicator of the participants (dis)comfort. Particularly, younger participants were more likely to be uncomfortable with their videos recorded when compared to older participants.
@@
While this reluctance indicates distrust in the system, a conclusion about H4 cannot be reached. We speculate that this distrust is largely due to the conjectural nature of the questions. The general population is conscious of their data privacy and hence distrustful of companies (mis)using their data. The questions asked in this context were phrased in a manner that did not specify what the collected recordings would be used for and how that data would be handled. This ambiguous phrasing could have led to participants conflating recent trends in data privacy and their concerns with the intended purpose of the question. These results along with the results shown in Figure 4 indicate that while drivers would prefer such a system, they are unwilling to accommodate its easiest plausible implementation. As such, we cannot offer a conclusion regarding H4.
VI Conclusion and Future Work
@@
In this paper, we present a user study that aims at testing driver preferences for the modes of communication between drivers and their assistant systems. We conducted a online user study (N=147) to test our four hypotheses regarding how drivers prefer their assistant systems communicate with them. Our user study involved participants viewing a set of videos and answering questionnaires about the videos, as well as questionnaires about an adaptive communication module. We validate our hypotheses that driver preferences for the modes of communication depend on the critical nature of the driving scenario and their attention state. We also affirm that drivers would feel safer with and hence prefer an adaptive communication module in their vehicles. Lastly, we hypothesised that drivers would trust vehicles that autonomously adapted their mode of communication, but found inconclusive evidence for confirmation.
VI-A Limitations
@@
While our work can provide automotive designers with important guidelines for ADAS development, it has the following limitations. Firstly, our study was conducted online where researchers had no interaction with participants and hence had to rely on metrics such as completion time and validation questions to verify responses. Additionally, researchers were unable to control the participants environment during the study and hence could not ensure minimal distractions. Though the videos shown were made of the highest possible resolution, compression on the hosting website as well as the quality of the participants internet connection at the time of completing the survey could have severely hindered the quality of the videos. This reduction in quality could lead to participants missing key details in the videos that could impact their responses. In general, even though we gained important insights into driver behavior through the study, we cannot provide concrete recommendations without further naturalistic studies, which we leave for future work. Secondly, two of the four hypotheses tested relied on participants understanding of a hypothetical system that they have no prior experience with. Therefore, any extrapolations made from the results presented in this paper should be thoroughly verified and tested independently. Lastly, the videos shown as part of the survey were recorded in a driving simulator and hence can only represent real world driving up to a reasonable level of accuracy.
VI-B Future Work
@@
Though the work presented here has the above limitations, the results discussed lead to following interesting avenues for future work. The results for H1 and H2 provide a basis for vehicle manufacturers to revisit traditional modes of communication for critical information in vehicles. In particular, the blind spot icon on mirrors is currently a popular mode for presenting information about the vehicles blind spot [10]. The results presented in this work show that other modes of communication or combinations of multiple modes might be preferred for that purpose. Conversely, modes of communication that rely on non-flashing icons and natural language text on the dashboard are not popular choices for most scenarios tested. However, recent designs for vehicles show an increase in the use of such modes [41]. Additional in-car studies to test the utility of these modes would provide further clarity into the most efficient ways of communicating important information to drivers. Finally, the development and testing of an adaptive communication module would greatly help extend the frontier of driver-automation communication due to the confirmation of H3. These adaptive systems for communication will also aid in providing a more comfortable driving experience and improve overall road safety.
Acknowledgements

### v2

@@ 3.3. Adaptive Communication Module Questionnaire
We posit that drivers’ perceived sense of safety indicates a preference for a system as we assume that most rational drivers value safety. The dependent variable was how safe participants felt, and the independent variables were the demographic details of the participants [^3] Discussed in Appendix A.1..
Implementing an Adaptive Communication Module would require in-car driver monitoring to assess drivers’ current distraction state. Currently, gaze-based distraction estimation is a state-of-the-art non-invasive method to estimate driver distraction which only requires eye gaze data through the use of ubiquitous low-cost RGBD cameras (Kashevnik et al., 2021; Shen et al., 2022; Soleimanloo et al., 2019). While other methods of estimating driver distraction such as EEGs and heart rate monitors are more accurate and have been used in previous naturalistic user studies, they are more invasive and require additional complicated hardware (McDonnell et al., 2021; Tejero and Roca, 2021; Arakawa, 2021). Additionally, the general population is also unfamiliar with these devices and their inclusion would require additional context which would make the survey more complex and arduous. Therefore, participants were asked if they would be comfortable with their audio and facial videos being recorded by their vehicle. The dependent variable was how comfortable participants felt, and the independent variables were the demographic details of the participants. We verify this claim by asking participants about their distractions in the driving experience questionnaire [^4] Results are presented in the Appendix A.2.. We posit that the user’s comfort with a plausible detection system in addition to their perceived sense of safety indicates their trust in such a system,
3.4. Participants
Participants were recruited through Amazon MTurk, public social media channels such as Reddit and LinkedIn, and email. The use of these channels allowed for a wider variance in participant experiences and backgrounds. Of the 147 participants, 90 participants were recruited through Amazon MTurk. Only participants from Amazon MTurk were compensated with \$12 per hour. Three participants were excluded from the study as they did not meet the inclusion criteria. Therefore, a total of 144 participants (mean age=37; median age=39; range=18-65; 87 males; 57 females) were considered in the analysis presented in Section 4.
@@ 4. Results and Discussion
In this section, we analyze the responses obtained from our user survey. We derive conclusions regarding the usage of modes based on the driving scenario and recommend design changes for the communication paradigms between drivers and ADAS.
4.1. The relation between the modes and characteristics of a driving scenario
@@
Figure 3. The distribution of the modes of communication based on the hazardous nature of the driving scenario (left to right in decreasing order), regardless of distraction status. Each section of the vertical bars shows the number of times the mode was picked as the first (bottom), second (middle), and third (top) choice respectively.
[…]
[A stacked bar plot showing a change in preference for the mode of communication based on the hazardous nature of the driving task with repeated sound alerts being preferred for high risk events and spoken language alerts preferred for low risk events.]A stacked bar plot showing the frequency with which the different modes of communication were selected by participants for the different driving scenarios. The plot shows a change in preference for the mode of communication based on the hazardous nature of the driving task with repeated sound alerts being preferred for high risk events and spoken language alerts preferred for low risk events.
@@
As seen in Figure 3, participants prefer different modes of communication depending on the hazardous nature of the driving scenario. Particularly, participants prefer direct modes that can be interpreted quickly such as repeated sound alerts and vibrations for emergent, dangerous scenarios. For low-risk scenarios, participants prefer modes that serve as reminders and can be interpreted without haste, such as a singular sound beep or spoken language alerts. These conclusions are in line with the guidelines presented by Campbell et al. (2018) and the results presented by Du et al. (2021).
@@
Meanwhile, some modes such as non-flashing icons and text or images on the dashboard are not preferred by participants in any setting. These modes are consistently the least selected modes across all driving scenarios presented in the study. Notably, these modes are usually reserved for long term information regarding vehicle health and might be unsuitable for conveying information for the driving scenarios depicted in the study. However, a growing number of vehicles implement text and image alerts on the dashboard and central consoles as common modes for all types of automotive information (Tesla, Inc., 2023). We suggest that these modes should be used sparsely and with careful consideration.
@@
The difference between the frequency of the modes for the most and least hazardous scenario: Pedestrian collisions and Standing still at a Green light, clearly indicates that drivers prefer different modes of communication based on the risk involved in a scenario. This conclusion is also reaffirmed by the steady change of the frequency of modes as the risk in the scenario decreases, as seen from left to right in Figure 3. However, we note that the final rankings of the different modes are consistent across the first four driving scenarios. While this consistency indicates an overall unchanging sentiment, we note that the deviations in the mode frequencies for each scenario are significant as proven by the chi-squared tests. Moreover, there are significant deviations in the frequency of selections for each mode that imply the importance of personalization for an adaptive communication framework.
4.2. Evolution of driver preferences with changing attention
@@
The responses from the video questionnaires can be treated as a ranking of the top-3 preferences out of the 10 choices presented in Table 2. These top-3 rankings can be transformed to fully ranked lists of all the modes by allowing for ties. We choose to assign all modes that were not selected by users a tied rank of 4. As mentioned in Section 3.2, we test RQ2 by comparing the ranked lists for the non-augmented and augmented versions of the videos. We use the following rank correlation metrics to compare the responses for each video pair:
  1. (1)
@@
    Kendall’s $\tau$ (Kendall, 1938): The difference between the fraction of ranking pairs that appear in order and the fraction of ranking pairs that appear out of order. Values close to 1 indicate strong agreement while values close to -1 indicate strong disagreement.
  2. (2)
@@
    Spearman’s $\rho$ (Spearman, 1904): The Pearson correlation coefficient between the rank variables. Values close to 1 indicate agreement while values close to -1 indicate disagreement.
  3. (3)
@@
    Rank Biased Overlap (RBO) (Webber et al., 2010): The expected average overlap between the two rankings at incremental overlap depths. Values close to 1 indicate agreement, while values close to 0 indicates disagreement.
@@
The different metrics measure different types of correlation between the two rankings. Both Kendall’s $\tau$ and Spearman’s $\rho$ are unweighted measures and only check to see the ordering of the two rankings is the same without placing any weight on the change in ranking (i.e. scores for items with rank 1 and 2 would be the same as the score for items with rank 2 and 3). Whereas the RBO metric is more sensitive to changes in ranking as it assigns significance to the actual rank changes and thus weighs the changes in ranking more than just their order. All analysis was carried out in python using the scipy and rbo packages (Virtanen et al., 2020; Chen et al., 2021).
@@
Table 3. The Correlations Between the Ranked Modes of Communication for the Non-augmented and Augmented Videos. The p-values for all metrics were less than 0.0004.
@@
Table 3 shows the correlations for these metrics aggregated according to the driving scenario. The correlations were aggregated as simple averages, and the p-values were aggregated using the Fisher exact test (Fisher, 1922). All metrics show that the rankings for the non-augmented and augmented versions are not fully correlated but only slightly positively correlated. We speculate that this result is a consequence of the modes of communication being more closely related to the critical nature of the driving scenario than the current distraction state. We further conjecture that the positive correlation implies that while driver preferences for the mode of communication change, drivers are not completely changing their minds on what modes best suit the current driving scenario. We also observe that the correlation between the rankings decreases as the scenarios become less hazardous. The correlation values ranging from 0.5 to 0.7 reinforces this belief.
Additionally, we believe that these values are also a result of converting the partially ranked top-3 list to a fully ranked list with ties at rank 4. This policy ensures that a large portion of the ranking lists are the same, since 7 out of 10 options were all ranked at the same rank. This procedure heavily biases the correlation observed between the rankings to be positive.
4.3. Sentiments towards an Adaptive Communication Module
@@
Figure 4. The distribution of responses for perceived safety in a car with an AdaCoM Module as a driver (top) and passenger (bottom).
[…]
[Two waffle plots showing participant responses to their perceived safety in a vehicle with AdaCoM that shows that over 77% of participants would feel safer with AdaCoM.]Two waffle plots showing participant responses to their perceived safety in a vehicle with an Adaptive Communication Module (AdaCoM) that shows that over 80% of participants would feel safer as drivers and over 77% of participants would feel safer as passengers.
@@
Figure 5. The distribution of responses for their comfort with audio (top) and visual (bottom) recordings in their vehicles.
[…]
[Two waffle plots showing participant responses to their comfort in being recorded by their vehicle that shows that over 60% of participants would feel uncomfortable the recording.]Two waffle plots showing participant responses to their comfort in being recorded by their vehicle that shows that over 60% of participants would feel uncomfortable with their audio being recorded and 67% of participants would feel uncomfortable with their video being recorded.
@@
The responses for participants perceived safety in a vehicle with an adaptive communication module are shown in Figure 5. The mean score (1=Strongly disagree; 5=Strongly agree) for perceived safety as a driver and passenger are $3.93\pm 0.91$ and $3.95\pm 1.03$, respectively, i.e.  over 80% and 77% of respondents believe that they would feel safer in a vehicle with AdaCoM as a driver and as a passenger, respectively.
@@
The responses for the participants comfort with a plausible implementation of an AdaCoM system are shown in Figure 5. The mean scores for comfort with audio and video recordings are $2.32\pm 1.20$ and $2.14\pm 1.19$, respectively. Over 60% and 67% of respondents are uncomfortable having their audio and video recorded by their vehicles, respectively.
@@
We speculate that this distrust is largely due to the conjectural nature of the questions asked to the participants. The general population is conscious of their data privacy and hence distrustful of companies (mis)using their data. The questions asked in this context were phrased in a manner that did not specify what the collected recordings would be used for and how that data would be handled. This ambiguous phrasing could have led to participants conflating recent trends in data privacy and their concerns with the intended purpose of the question. These results along with the results shown in Figure 5 indicate that while drivers would prefer such a system, they are unwilling to accommodate its easiest plausible implementation.
5. Conclusion and Future Work
@@
In this paper, we present a user study that aims at testing driver preferences for the modes of communication between drivers and their assistant systems. Our survey involved participants viewing a set of videos and answering questionnaires about the videos, as well as questionnaires about an adaptive communication module. We confirm that driver preferences for the modes of communication depend on the critical nature of the driving scenario and their attention state. We also affirm that drivers would feel safer with and hence prefer an adaptive communication module in their vehicles.
5.1. Limitations
@@
While our work can provide automotive designers with important guidelines for ADAS development, it has the following limitations. Firstly, our study was conducted online where we had no interaction with participants and hence had to rely on metrics such as completion time and validation questions to verify responses. Additionally, we were unable to control the participants environment during the study and hence could not ensure minimal distractions, beyond participant assurance. Though the videos shown were made of the highest possible resolution, compression on the hosting website as well as the quality of the participants internet connection at the time of completing the survey could have severely hindered the quality of the videos. This reduction in quality could lead to participants missing key details in the videos that could impact their responses. In general, even though we gained important insights into driver behavior through the study, we cannot provide concrete recommendations without further naturalistic studies, which we leave for future work. Therefore, any extrapolations made from the results presented in this paper should be thoroughly verified and tested independently. Lastly, the videos shown as part of the survey were recorded in a driving simulator and hence can only represent real world driving up to a reasonable level of accuracy.
5.2. Future Work
@@
Though the work presented here has the above limitations, the results discussed lead to the following interesting avenues for future work. In particular, the blind spot icon on mirrors is currently a popular mode for presenting information about the vehicles blind spot (Campbell et al., 2018). The results presented in this work show that other modes of communication or combinations of multiple modes might be preferred for that purpose. Conversely, modes of communication that rely on non-flashing icons and natural language text on the dashboard are not popular choices for most scenarios tested. However, recent designs for vehicles show an increase in the use of such modes (Tesla, Inc., 2023). Additional in-car studies to test the utility of these modes would provide further clarity into the most efficient ways of communicating important information to drivers. Finally, the development and testing of an adaptive communication module would greatly help extend the frontier of driver-automation communication.
Acknowledgements.